OpenStudy (mhchen):

What is the least positive integer 'n' such that the product of 2016 and 'n' ends in the digits 00?

OpenStudy (518nad):

2016 = 4*504 4*2*252 4*4*126 4*4* much work.. 2^5*3^2*7

OpenStudy (518nad):

so we gotta get a 100 in there

OpenStudy (518nad):

2^5*3^2*7 9*7,4*4*2 so... its probably 25?

OpenStudy (mhchen):

Hmmm.... 5 * 6 = 30 .........2.....You are right :o

OpenStudy (518nad):

ya well

OpenStudy (518nad):

100 = 10^2=2^2*5^2 sooo and we got 2016 = 2^5*3^2*7 so we are missing the 5^2 in there to get to a 100 multiple

OpenStudy (518nad):

moree ask moree

OpenStudy (518nad):

to justify it more clearly incase they want like proper solution 2016 = 2^2*(2^3*3^2*7) the part in the bracket has no factor 10 so it doesnt end in 0 now if we multiple any number not ending in 0 by 100 we will get xxxx00 so we gotta multiply by 25 2^2*(2^3*3^2*7) --> (25*2^2)*(2^3*3^2*7) now its in form 100*some number not ending in 0

OpenStudy (mhchen):

Yeah that makes sense. Factorizationmunizationishunnnn and getting products 2*5 = 10 * 10 blah blah blah.

OpenStudy (518nad):

2*5 = 10 not 10*10

OpenStudy (518nad):

:)